deathscarewolf:

Eric plotted the graph below to show the relationship between the temperature of his city and the number of cups of lemonade he sold daily: (in comments) Part A: Describe the relationship between the temperature of the city and the number of cups of lemonade sold. Part B: Describe how you can make the line of best fit. Write the approximate slope and y-intercept of the line of best fit. Show your work, including the points that you use to calculate the slope and y-intercept.

deathscarewolf:

snowflake0531:

Let's first do Part A, Describe the relationship between the temperature of the city and the number of cups of lemonade sold. this should be obvious, seeing as when the temperature grows, the amount of lemonade sold also grows

deathscarewolf:

ya i could see that

snowflake0531:

Okay, for line of best fit, you can see that the line is kinda right there, going through the points (30,4)and (90,2) Can you figure out the slope?

deathscarewolf:

(50,2)?

snowflake0531:

The slope, do you know the slope thingy? how to find it between 2 points?

deathscarewolf:

kinda forgot

AZ:

\(\Large\text{Slope} = \dfrac{\color{green}{y_2} - \color{orange}{y_1}}{\color{cyan}{x_2}-\color{red}{x_1}}\) where you have two points \(\Large (\color{red}{x_1}, \color{orange}{y_1})\) and \(\Large (\color{cyan}{x_2}, \color{green}{y_2})\) So your two points are \((\color{red}{30}, \color{orange}{4})\) and \((\color{cyan}{90}, \color{green}{20})\)

snowflake0531:

So can you use ^^^^^^^ and find the slope between teh two points?

deathscarewolf:

idk to be honest xD

snowflake0531:

We do (20-4)/(90-30) can you simplify that?

deathscarewolf:

(16,60)?

snowflake0531:

The slope isn't a point, it's a number In this case, 16/60, which is simplified to 4/15

deathscarewolf:

before i did both and got 0.26667

snowflake0531:

YOu want exact form, so you don't convert it into a fraction, it stays at 4/15

deathscarewolf:

ok

snowflake0531:

So now we want to figure out the y-intercept So we take the slope and a coordinate point 4=4/15(30) + b 4 = 8+b b=-4

snowflake0531:

So, we have \[y=\frac{ 4 }{ 15 } x -4\]

deathscarewolf:

ok

snowflake0531:

so there's your answer.-.

deathscarewolf:

ty

snowflake0531:

yw..